OpenStudy (anonymous):

Solve: 2sqrt{x+5}+3sqrt{x-3}=9

OpenStudy (anonymous):

I keep coming up with 88/13

OpenStudy (zarkon):

4

OpenStudy (anonymous):

4 is the correct answer. to solve it 1.) first square on both sides and simplify. You will get \[12\sqrt{(x+5)(x-3)} = 88-13x\] 2.) then again square it on both the sides and simplify. You will end up with a quadratic equation. 3.) If you solve that quadratic equation you will get x=4 as one of the solutions. 4.) You can verify it by substituting it back into the equation given.

OpenStudy (anonymous):

if i square both sides in the first step doesn't that make the equation \[4(x+5)+9(x-3)=81\]

OpenStudy (zarkon):

no

OpenStudy (zarkon):

\[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

o right thanks.

OpenStudy (anonymous):

\[2\sqrt{x+5}+3\sqrt{x-3}=9\] that is the equation you were using the first time right?

OpenStudy (anonymous):

\[(2\sqrt{x+5}+3\sqrt{x-3})^{2}=9^{2}\] \[(2\sqrt{x+5})^{2}+2*(2\sqrt{x+5})*(3\sqrt{x-3})+(3\sqrt{x-3})^{2}=9^{2}\]

OpenStudy (anonymous):

\[4(x+5)+12*(\sqrt{x+5})*(\sqrt{x-3})+9(x-3) = 81\]

OpenStudy (anonymous):

if i square this equation i get 144(x+5)(x-3)=7744-2288-169x^2

OpenStudy (anonymous):

For (2sqrt(x+5) + 3sqrt(x-3))^2 I come up with 13x-7+12sqrt(x+5)(x-3)

myininaya (myininaya):

i got something ugly so far 25x^2-2576x+9904=0

OpenStudy (akshay_budhkar):

on simplifying you get 88-13x=12√(x+5)(x-3) right?

OpenStudy (akshay_budhkar):

then square it..

OpenStudy (akshay_budhkar):

but it looks ugly then

OpenStudy (anonymous):

apparently the answer is 4. I typed into wolframs and that what it gave me

OpenStudy (akshay_budhkar):

@myininaya you squared that???

OpenStudy (anonymous):

if you square 88-13x. I get 7744-2288=169x^(2)

OpenStudy (akshay_budhkar):

144(x+5)(x-3)=7744-2288-169x^2 do we get this... man this is very ugly

myininaya (myininaya):

i have to check my answers now since i raised both sides to even power

myininaya (myininaya):

but like 99.04 is totally not going to work

myininaya (myininaya):

so the answer is just 4

myininaya (myininaya):

can you read my attachment raioner?

OpenStudy (anonymous):

yes im looking it over now

OpenStudy (anonymous):

trying to find out where i went wrong thanks for doing that for me.

myininaya (myininaya):

np those were some ugly numbers

OpenStudy (akshay_budhkar):

ya right.. but why is 99.04 wrong? because of the domain of x right?

myininaya (myininaya):

whell if you plug it in both sides are not the same

OpenStudy (anonymous):

this is part of my take home test and and the entire test is like this.

OpenStudy (akshay_budhkar):

root of something is geater than 0.. so x-3 greater than 0 and x=5 greater than 0...

OpenStudy (akshay_budhkar):

*x+5

OpenStudy (anonymous):

Ok, when you square both sides of the equation, you will come up with (x-4)(25x-2476)=0 so you'll want to set both sides of the equation to equal 0 x-4=0, 25x-2476=0 so you'll have x=4 and x=2476/25 now plug into the original equation start with 4 (because it looks easy :) ) and you'll find 4 works, you can test 2476/25 as well but no need, myininaya already confirmed x=4 :D

OpenStudy (akshay_budhkar):

ya but you can also take the domain stuff in consideration no need to verify then..

myininaya (myininaya):

whenever you raise both sides to an even power you should always check your answers

OpenStudy (akshay_budhkar):

ya squaring increasing the number roots..

OpenStudy (anonymous):

Very true

OpenStudy (akshay_budhkar):

he will get irritated finding the answer and verifying 99.04 will surely eat his head up.. lol!

myininaya (myininaya):

i would make my calculator do the arithmetic which i did lol

myininaya (myininaya):

gn guys